OpenStudy (anonymous):

Find the sum of the arithmetic sequence. -4, -1, 2, 5, 8, 11, 14

OpenStudy (anonymous):

hey it is simple.. we know that sum is given by =n/2*(2a+(n-1)*d) here n=7, a=-4 and d=3 putting this together we get sum=7/2*(2*-4+(7-1)*3) =35

OpenStudy (anonymous):

@irshadno1 thank you